A would takes 6 hours to paint a wall working alone. B takes 4 hours. C 3 hours
A starts painting at noon, at 1.00 pm B joins him, C at 1.30 pm. At what time do they finish the painting ?

Answer by Edwin McCravy(20060)

A would takes 6 hours to paint a wall working alone. B takes 4 hours. C 3 hours
A starts painting at noon, at 1.00 pm B joins him, C at 1.30 pm. At what time do they finish the painting ?
Let t = the time in hours after 1:30PM until the painting is finished.

Make this chart and fill in t for the time after 1:30PM until the painting
is done.

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |                |      |
B working alone     |                |      |
C working alone     |                |      | 
A from 12 to 1PM    |                |      |
A&B from 1 to 1:30PM|                |      | 
A&B&C after 1:30PM  |                |  t   | 


>>...A would takes 6 hours to paint a wall working alone.
B takes 4 hours. C 3 hours...<<
Fill in the times for A, B, and C and 1 for the number of walls painted
in those times.

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |
B working alone     |       1        |  4   |
C working alone     |       1        |  3   | 
A from 12 to 1PM    |                |      |
A&B from 1 to 1:30PM|                |      | 
A&B&C after 1:30PM  |                |  t   | 

Next fill in the Walls/hour rates for A, B and C individually, by 
dividing their number of walls painted, which is 1, by their times
in hours to paint 1 wall working alone:

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |                |      |
A&B from 1 to 1:30PM|                |      | 
A&B&C after 1:30PM  |                |  t   | 

Fill in A's rate of 1/6 walls/hour also when he paints from 12 noon to 1PM
and 1 hour for the time:

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |                |  1   |        1/6 
A&B from 1 to 1:30PM|                |      | 
A&B&C after 1:30PM  |                |  t   | 

Add A's rate and B's rate to find their combined rate:
1/6 + 1/4 = 2/12 + 3/12 = 5/12 and 1/2 hour for their time working
together from 1PM to 1:30PM

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |                |  1   |        1/6 
A&B from 1 to 1:30PM|                | 1/2  |       5/12  
A&B&C after 1:30PM  |                |  t   | 

Add A's, B's and C's rate to find the combined rate for all three
after 1:30PM

1/6 + 1/4 + 1/3 = 2/12 + 3/12 + 4/12 = 9/12 = 3/4

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |                |  1   |        1/6 
A&B from 1 to 1:30PM|                | 1/2  |       5/12  
A&B&C after 1:30PM  |                |  t   |        3/4

Fill in the fraction of a wall painted by A alone from 12 noon to 1PM
by multiplying his rate times his time: (1/6)(1) = 1/6

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |      1/6       |  1   |        1/6 
A&B from 1 to 1:30PM|                | 1/2  |       5/12  
A&B&C after 1:30PM  |                |  t   |        3/4

Fill in the fraction of a wall painted by A&B from 1PM to 1:30PM
by multiplying their combined rate times their time: (1/2)(5/12) = 5/24

                      Walls painted  | Time | Rate in Walls/hour
A working alone     |       1        |  6   |        1/6
B working alone     |       1        |  4   |        1/4
C working alone     |       1        |  3   |        1/3
A from 12 to 1PM    |      1/6       |  1   |        1/6 
A&B from 1 to 1:30PM|     5/24       | 1/2  |       5/12  
A&B&C after 1:30PM  |   (3/4)t       |  t   |        3/4

Fill in the fraction of a wall painted by A&B&C after 1:30PM
by multiplying their combined rate times their time: (3/4)t 

    The equation is obtained by adding the three fractions
    of a wall painted beginning at 12 noon, and setting that
    sum equal to 1 wall:

            1/6 + 5/24 + (3/4)t = 1

Clear of fractions by multiplying through by 24:

                    4 + 5 + 18t = 24
                        9 + 18t = 24
                            18t = 15
                              t = 15/18
                              t = 5/6 of an hour

5/6 of an hour is (5/6)(60) = 50 minutes.  So the painting was 
finished at 50 minutes after 1:30PM or 2:20PM.

Edwin